Equivalent Units and Related Costs; Cost of Production Report; Entries

Dover Chemical Company manufactures specialty chemicals by a series of three processes, all materials being introduced in the Distilling Department. From the Distilling Department, the materials pass through the Reaction and Filling departments, emerging as finished chemicals.

The balance in the account Work in Process—Filling was as follows on January 1:

Work in Process—Filling Department
(1,900 units, 40% completed):
Direct materials (1,900 x $16.1) $30,590
Conversion (1,900 x 40% x $10.5) 7,980
$38,570

The following costs were charged to Work in Process—Filling during January:

Direct materials transferred from Reaction
Department: 24,500 units at $15.9 a unit $389,550
Direct labor 140,070
Factory overhead 134,574

During January, 24,300 units of specialty chemicals were completed. Work in Process—Filling Department on January 31 was 2,100 units, 90% completed.

Required:

1. Prepare a cost of production report for the Filling Department for January. If an amount is zero, enter "0". If required, round your cost per equivalent unit answers to two decimal places.

Dover Chemical Company
Cost of Production Report-Filling Department
For the Month Ended January 31
Unit Information
Units charged to production:
Inventory in process, January 1
Received from Reaction Department
Total units accounted for by the Filling Department
Units to be assigned costs:
Equivalent Units
Whole Units Direct Materials Conversion
Inventory in process, January 1
Started and completed in January
Transferred to finished goods in January
Inventory in process, January 31
Total units to be assigned costs
Cost Information
Cost per equivalent unit:
Direct Materials Conversion
Total costs for January in Filling Department $ $
Total equivalent units
Cost per equivalent unit $ $
Costs assigned to production:
Direct Materials Conversion Total
Inventory in process, January 1 $
Costs incurred in January
Total costs accounted for by the Filling Department $
Costs allocated to completed and partially completed units:
Inventory in process, January 1 balance $
To complete inventory in process, January 1 $
Cost of completed January 1 work in process $
Started and completed in January $
Transferred to finished goods in January $
Inventory in process, January 31
Total costs assigned by the Filling Department $

2. Journalize the entries for (1) costs transferred from Reaction to Filling and (2) the cost transferred from Filling to Finished Goods. If an amount box does not require an entry, leave it blank.

(1)
(2)

3. Determine the increase or decrease in the cost per equivalent unit from December to January for direct materials and conversion costs. If required, round your answers to two decimal places.

Increase or Decrease Amount
Change in direct materials cost per equivalent unit $
Change in conversion cost per equivalent unit $

4. Discuss the uses of the cost of production report and the results of part (3).

The cost of production report may be used as the basis for allocating product costs between   and  . The report can also be used to control costs by holding each department head responsible for the units entering production and the costs incurred in the department. Any differences in unit product costs from one month to another, such as those in part (3), can be studied carefully and any significant differences investigated.
Answer #1

Part 1

DOVER CHEMICAL COMPANY

Cost of Production Report—Filling Department

For the Month Ended January 31

Unit information

Units charged to production:

Inventory in process, January 1

1900

Received from Reaction Department

24500

Total units accounted for by the Filling Department

26400

Units to be assigned costs:

Equivalent units

Whole units

Direct materials

Conversion

Inventory in process, January 1

1900

0

1140

Started and completed in January

22400

22400

22400

Transferred to finished goods in January

24300

22400

23540

Inventory in process, January 31

2100

2100

1890

Total units to be assigned cost

26400

24500

25430

1900*(1-40%) =1960

24300-1900 = 33000

2100*90% = 1890

Cost information

Costs per equivalent unit:

Direct materials

Conversion

Total costs for January in Filling Department

389550

274644

Total equivalent units

24500

25430

Cost per equivalent unit

15.90

10.80

Costs charged to production:

Direct materials

Conversion

Total

Inventory in process, January 1

38570

Costs incurred in January

664194

Total costs accounted for by the Filling Department

702764

Cost allocated to completed and partially completed units:

Inventory in process, January 1 balance

38570

To complete inventory in process, January 1

0

12312

12312

Cost of completed January 1 work in process

50882

Started and completed in January

356160

241920

598080

Transferred to finished goods in January

648962

Inventory in process, January 31

33390

20412

53802

Total costs assigned by the Filling Department

702764

1140*10.80 = 12312

22400*15.90 =356160

22400*10.80 =241920

2100*15.90 =33390

1890*10.80 = 20412

Part 2

No.

Account titles and explanation

Debit

Credit

1

Work in Process—Filling Department

389550

Work in Process—Reaction Department

389550

2

Finished Goods

648962

Work in Process—Filling Department

648962

Part 3

Change in Direct materials cost per equivalent unit

Decrease

$0.20

Change in Conversion cost per equivalent unit

Increase

$0.30

16.10-15.90 = 0.20

10.50-10.80 = 0.20

Part 4

The cost of production report may be used as the basis for allocating product costs between Work in Process and Finished Goods. The report can also be used to control costs by holding each department head responsible for the units entering production and the costs incurred in the department. Any differences in unit product costs from one month to another, such as those in part (3), can be studied carefully and any significant differences investigated.